A new still unsharpened, wooden pencil has the form of a cylinder. It is 19cm long and has a diameter of 0.8cm.The cylinder hole for the core in the middle of the pencil has a diameter of 0.2cm
1) calculate how much cm3 of wood is needed for such a pencil

Answer by mananth(16949)

outer volume - inner volume = volume of wood for the pencil
outer radius =0.8/2 = 0.4 cm
length = 9 cm
inner radius = 0.2/2 = 0.1
length = 9 cmm
V cylinder = pir^2h
V(whole pencil) = pi * 0.4^2*9
V (core) = pi*0.1^2*9
V ( wood) = pi*0.4*2*9-pi*0.1^2*9
= 9*pi(0.4^2-0.1^2)
=4.24 cm^3